# Helfrich cylinder Alexander Meiners, Hannes Uecker

### v1

@@ Abstract
We study bifurcations from the cylinder governed by the Helfrich equation. The bifurcating solutions are in three classes, axisymmetric, non axisymmetric without Gaussian curvature and completely non axisymmetric solutions. Depending on the spontaneous curvature we find stability regimes of the cylinder and destabilizing bifurcations to spatially periodic solutions. To get a complete picture we use well established local bifurcation methods. Numerical continuation provides global stability statements of the bifurcations.
@@ 1 Introduction
We consider the Helfrich equation
$\dot{X}=\left(\Delta H+2H(H^{2}-K)+2c_{0}K-2(c_{0}^{2}+\lambda_{1})H-\lambda_{2}\right)N,$ (1)
with $c_{0}$ the spontaneous (or preferred curvature) and $\Lambda:=(\lambda_{1},\lambda_{2})$ are Lagrange multipliers for the area and the volume, respectively. The cylinder $\mathcal{C}_{r}$ of radius $r$ is a stationary solution to (1) iff
@@
Typically for compact manifolds the integral
$E(X,c_{0})=\int_{X}a(H-c_{0})^{2}+bK\,{\rm d}S$ (3)
is called the Helfrich energy. The area of such manifold is define by
${\cal A}(X):=\int_{X}\,{\rm d}S$ (4)
and for cylindrical surfaces the volume is
${\cal V}(X)=\frac{1}{2}\int_{X}\frac{(r+u)^{2}}{g^{1/2}}\,{\rm d}S=\frac{1}{2}\int_{I\times\mathbb{S}^{1}}(r+u)^{2}\,{\rm d}(x,\varphi).$ (5)
First introduced in [Hel73], for compact manifolds an area and volume constraint was introduced, leading to the Lagrangian
$F(X,c_{0})=E(X,c_{0})+\lambda_{1}({\cal A}(X)-{\cal A}_{0})+\lambda_{2}({\cal V}(X)-{\cal V}_{0}),$ (6)
with Euler-Lagrange equation (1) with $a=1$ and $b=0$, for variation in normal direction of $X$. In the case with boundary additional boundary terms come up, defining natural boundary conditions, [PP22]. But in the case of $\mathcal{C}_{r}$ the energy is not well defined, rather on each compact portion of length $L$, call this $\mathcal{C}_{r}(L)$. However on $\mathcal{C}_{r}(L)$ we have a scaling, that for $E(\mathcal{C}_{\alpha r}(L),c_{0})=E(\mathcal{C}_{r}(\alpha L),c_{0}/\alpha)$. So near a compact cylinder, it is sufficient to consider $r=1$ with a variable length and abbreviate $\mathcal{C}=\mathcal{C}_{1}$.
Associated to (6) is a $L^{2}$-gradient flow of the form
$\begin{pmatrix}V_{t}:=\left\langle\dot{X},N(t)\right\rangle\\ 0\end{pmatrix}=\begin{pmatrix}G(X(t),\Lambda)\\ q(X(t))\end{pmatrix},\quad q(X(t))=\begin{pmatrix}{\cal A}(X(t))-{\cal A}_{0}\\ {\cal V}(X(t))-{\cal V}_{0}\end{pmatrix}$ (7)
where $G(X(t),\Lambda)$ is the rhs of (1) and $\lambda_{1}(t)$ and $\lambda_{2}(t)$ are time depending functions, such that the constraints for ${\cal A}(X)$ and ${\cal V}(X)$ are fulfilled. We will prove well-posedness result of (7) for initial conditions $X_{0}$ close to $\mathcal{C}_{r}$ + global existence.
1.1 Introductory example
For our first observation we use the continuation package pde2path, especially with the functions of the Xcont library. For short review of the numerical procedure see subsection A.1 and for a complete introduction see [MU24]. In our first experiments we try to capture the influence of the bifurcation wrt $c_{0}$. We use periodic boundary conditions, where the period sufficiently large ($L=10$) to detect primary bifurcation in the periodic boundary conditions. Neumann typ boundary conditions would also be a reasonable choice. The first test value is $c_{0}=0$ with the focus in the stability of the cylinder and destabilizing bifurcations. However an extended bifurcation diagram is discussed in section 5.
(a) (b)
Refer to caption Refer to caption Refer to caption Refer to caption
Figure 1: Bifurcation diagram (a) of (1) for the case $L=10$ and $c_{0}=0$. With sample solutions in (b) for the first bifurcation 0-BP1 in dark red to an axisymmetric branch. Then two branches bifurcating from 0-BP2 to the unstable spot like on the orange and stable spiral (strip like) solutions on the green branch.
For a given $r$ and $c_{0}$ equation (2) describes a family of active parameter vector $(\Lambda,V)$ on the trivial branch, with $V=V_{0}$. While the notion of stability will discussed rigorously in section 2, we shall comment on that now. We will consider the Helfrich flow projected on area and volume preservation. The parameters $\lambda_{i}$ are the Lagrange multipliers to the area and volume constraint. Hence in the dynamical problem (7) the $\lambda_{i}$’s are functions of $t$ to ensure the constraints. In the continuation the stability is calculated for a fixed primary continuation parameter, hence stability information is not valid if $\lambda_{i}$’s are primary. However we can easily change the parameter vector, e.g. to $(V,\Lambda)$ for the eigenvalue calculation and the detection of the bifurcation points, while volume change along the bifurcating branch ins allowed. Due to the scale invariance of (3) on the cylinder and the two constraints we introduce the reduced volume as a uniform parameter describing the phase diagram, for more details see section 5.
[…]
This point of view of allow the reduced volume to change along the branch, while considering volume preserving stability information is mainly motivated by the closed vesicle situation discussed for example in [JSR93]. Roughly s peaking the sphere is the unique optimum of maximal volume by minimal area. Hence volume and area preserving branches from the sphere do not exist. However the cylinder can be destabilized preserving the area and the volume while changing the langth noticed already in [ZcH89].
The first samples are done with $c_{0}=0$ presented in Figure 1. We find a narrow range of stability of the trivial branch 0, with two branch points at the end. At 0-BP1 (first number corresponds to the value of $c_{0}$) with $\lambda_{1}\approx?$ the bifurcating branch is axisymmetric. In contrast to the surface diffusion flow and considerations in [LS13] constant mean curvature surfaces of revolution are not solutions to (1) with $c_{0}=0$. For $c_{0}\neq 0$ this case can happen if $c_{0}=H(X)$. 0-BP1 has multiplicity two, corresponding to a translation of $L/4$ in $x$ direction and the bifurcation is a supercritical pitchfork, with the hidden symmetry that translation of $L/2$ corresponds to the other direction of the pitchfork.
The second destabilization 0-BP2 lead to a non-axisymmetric solution with axial wave number 1. However due to the symmetry of the kernel solutions 0-BP2 has multiplicity 4. with kernel function, rotated and translated. Due to the high multiplicity two branches bifurcate a corkscrews branch being stable and a wave like branch unstable.
1.2 Organization
First we shall set up the necessary differential geometry and rewrite(7) as an equation in the normal displacement of the cylinder. This gives a constrained gradient flow in $u$ associate with (3). In contrast to many others we establish maximal regularity of the parabolic problem in Sobolev spaces. This gives a short time of the flow. Due to the translation invariance of $\mathcal{C}_{r}$ close to the cylinder there is a two dimensional center manifold of cylinders with the same radius.
Again using the maximal regularity we can utilize the abstract result from Simonett und so that nonlinear stability follows from linearized stability. Since $\mathcal{C}_{r}$ has constant mean and Gauß curvature, the eigenvalues and eigenfunctions of the linearized stationary problem can be computed explicitly. Hence we derive the amplitude equations With this characterization we can prove the
@@ 2 Preliminaries and well-posedness of the flow
We introduce the basic differential geometric preliminaries in order to write (7) as a constrained parabolic quasi-linear equation in the normal displacement $u$ of the cylinder. This construction is as in [LS16], and hence we refrain from extensive computational details and focus on the structure and mapping properties of the resulting equations. From these we obtain maximal regularity results and hence local existence for (7). For general background on differential geometry we recommend [Tapp16, UY17].
@@ 2.1 Preliminaries
$\begin{array}[]{rl}\Phi:{\mathbb{R}}\times\T&\rightarrow{\mathbb{R}}^{3}\\ (x,\varphi)&\mapsto\begin{pmatrix}x,r\cos(\varphi),r\sin(\varphi)\end{pmatrix}^{T}\end{array}$ (10)
[…]
@@
assuming $u<r$, such that $X$ is still an embedding of $\T\T{\mathbb{R}}\times\T$; explicitly,
@@
[…]
For $u,v\in C^{\infty}_{\text{per}}({\cal C}_{L}(r))$ sufficiently small, the constraints (LABEL:c1) around $X(u)=\mathcal{C}_{r}-u\,N$ expand as ${\cal A}(u+v)={\cal A}(u)+\left\langle{\partial}_{u}{\cal A}(u),v\right\rangle+O(\|v\|^{2})$ with
$\left\langle{\partial}_{u}{\cal A}(u),v\right\rangle=2\int_{X}H(u)\eta(u)v\,{\rm d}S,$ (16)
and second derivative at ${\cal C}_{L}$
$\left\langle{\partial}_{u}^{2}{\cal A}(0)v,v\right\rangle=\int_{X}-(\Delta_{0}v+2(H^{2}(0)-K(0))v)v\,{\rm d}S$ (17)
and the volume
@@
expands as ${\cal V}(u+v)={\cal V}(u)+\left\langle{\partial}_{u}{\cal V}(u),v\right\rangle+\frac{1}{2}\left\langle{\partial}_{u}^{2}{\cal V}(u)v,v\right\rangle$ with
$\left\langle{\partial}_{u}{\cal V}(u),v\right\rangle=\int_{X}\eta(u)v\,{\rm d}S,\,\left\langle{\partial}_{u}^{2}{\cal V}(u)v,v\right\rangle=\int_{X}g^{-1/2}(u)v^{2}\,{\rm d}S.$ (19)
[…]
Since $\Lambda$ appears linearly in $G(u,\Lambda)$, the equation for the first derivative $\frac{\rm d}{{\rm d}t}q(u)=0$ gives a linear system of equations
$\mathcal{D}\Lambda=-\begin{pmatrix}\left\langle{\partial}_{u}E,{\partial}_{u}{\cal A}(u)\right\rangle\\ \left\langle{\partial}_{u}E,{\partial}_{u}{\cal V}(u)\right\rangle\end{pmatrix},$ (20)
@@
$\mathcal{D}:=\left\langle{\partial}_{u}q(u),{\partial}_{\Lambda}G(u,\lambda)\right\rangle=-\begin{pmatrix}4\int_{X}H^{2}(u)\eta^{2}(u)\,{\rm d}S&2\int_{X}H\eta^{2}(u)\,{\rm d}S\\ 2\int_{X}H\eta^{2}(u)\,{\rm d}S&\int_{X}\eta^{2}(u)\,{\rm d}S\end{pmatrix}.$ (21)
If $\det\mathcal{D}=4\left(\int_{X}\eta^{2}\,{\rm d}S\int_{X}H^{2}\eta^{2}\,{\rm d}S+\left(\int_{X}H\eta^{2}\,{\rm d}S\right)^{2}\right)\neq 0$, then the implicit functions theorem gives that in a neighborhood of $u$ there is a unique solution $\Lambda=\Lambda(u)$. Following [NY12], inserting $\Lambda(u)$ in (15a) gives a parabolic integro differential equation. However, if $u=0$, then $\det{\cal D}=0$, but the right hand side of (20) is orthogonal to $\ker{\cal D}$ (more generally, this holds at any constant mean curvature surface). Then a solution $\Lambda$ to (20) exists, and $\left\langle\Lambda_{\cal C},\ker{\cal D}^{T}\right\rangle=0$ gives that $2\lambda_{1}=\lambda_{2}$ and $\lambda_{1}=\frac{1}{3{\cal A}_{0}}\left\langle{\partial}_{u}E,1\right\rangle$, i.e., $\ker{\cal D}$ corresponds to the family of cylinders in (2); see [NY12] for further discussion (for the case of closed vesicles of spherical topology).

### v2

# Helfrich cylinders – instabilities, bifurcations and amplitude equations Alexander Meiners, Hannes Uecker

@@ Abstract
Combining local bifurcation analysis with numerical continuation and bifurcation methods we study bifurcations from cylindrical vesicles described by the Helfrich equation with volume and area constraints, with a prescribed periodicity along the cylindrical axis. The bifurcating solutions are in two main classes, axisymmetric (pearling), and non–axisymmetric (coiling, buckling, and wrinkling), and depending on the spontaneous curvature and the prescribed periodicity along the cylinder axis we obtain different stabilities of the bifurcating branches, and different secondary bifurcations.
Keywords. Spontaneous curvature, Helfrich energy, bifurcation analysis, numerical continuation
MSC 35B32, 58J55, 65P30
@@ 1 Introduction
The (dimensionless) Canham–Helfrich model [Can70, Hel73] (or spontaneous curvature model) for vesicles with lipid bilayer membranes is based on the energy
$E(X)=\int_{X}(H-c_{0})^{2}\,{\rm d}S,$ (1)
where $X$ is a (possibly time $t$ dependent) 2D manifold embedded in ${\mathbb{R}}^{3}$ with mean curvature $H$, and where the parameter $c_{0}\in{\mathbb{R}}$ models an energetically favorable spontaneous (or preferred) curvature; it depends on the properties of the lipid bilayer membrane and can be zero, or positive, or negative. Roughly speaking, the aim is to minimize $E$ for prescribed area ${\cal A}(X)$ and enclosed volume ${\cal V}(X)$, i.e., under the constraints
 $\text{ $q_{1}(X):={\cal A}(X)-{\cal A}_{0}=0$ and $q_{2}(X):={\cal V}(X)-{\cal V}_{0}=0$}.$ (2)
There are numerous works on this, for instance using the minimization of $E$ to explain the shape of red blood cells, see, e.g., [SZ89, SBL90, Pet83, SL95, Sei97]. [^1] For $c_{0}=0$, $E$ is also called Willmore energy. More generally, a term $b\int_{X}K\,{\rm d}S$ can be added to $E$, where the parameter $b\geq 0$ is called saddle–splay modulus, but due to the Gauss–Bonnet theorem $\int KdS$ is a topological invariant for closed $X$, and hence can be dropped from the minimization of $E$, and the same holds for the cylindrical topology considered here. The minimization proceeds by introducing the Lagrangian
$F(X,\Lambda)=E(X)+\lambda_{1}({\cal A}(X)-{\cal A}_{0})+\lambda_{2}({\cal V}(X)-{\cal V}_{0}),$ (3)
with Lagrange multipliers $\Lambda=(\lambda_{1},\lambda_{2})\in{\mathbb{R}}^{2}$, where $\lambda_{2}$ is the pressure difference between inside and outside; since we choose the inner normal, here negative $\lambda_{2}$ corresponds to lower pressure inside; $\lambda_{1}$ corresponds to a surface tension, and can also have either sign; see, e.g., [BHAVMC17] for discussion.
From (3) we obtain the Euler–Lagrange equation
$G(X,\Lambda):=\Delta H+2H(H^{2}-K)+2c_{0}K-2(c_{0}^{2}+\lambda_{1})H-\lambda_{2}=0$ (4)
under variations in normal direction together with the constraints (2), where $K$ is the Gaussian curvature, and $\Delta$ the Laplace–Beltrami operator on $X$. Many works also use (1) to study instabilities of tubular vesicles, e.g., [ZcH89, BZM94, FTGB+99, TSA+01], in particular the so–called pearling and coiling instabilities, see Fig.1(a,b), and Remark 1 for a review of the (physics) literature.
(a) (b)      (c)
[…]
In, e.g., [NY12, RSS24], the constrained $L^{2}$–gradient flow associated to (1) is studied, which can be written as
$\begin{pmatrix}V_{t}(t)\\ 0\end{pmatrix}=\begin{pmatrix}G(X(t),\Lambda(t))\\ q(X(t))\end{pmatrix}$, (5)
where $q=(q_{1},q_{2})$ and $V_{t}(t)$ is the normal velocity of $X$, and with now time dependent Lagrange multipliers $\Lambda(t)$ which are determined from (2), such that
$\frac{\rm d}{{\rm d}t}E(X(t),\Lambda(t))=-\|V(t)\|_{L^{2}(X(t))}^{2},\quad\frac{\rm d}{{\rm d}t}{\cal A}(X(t))=0,\quad\text{and }\frac{\rm d}{{\rm d}t}{\cal V}(X(t))=0,$ (6)
as long as the flow exists. We also refer to [Gar13] and [BDGP23] for introductions to, reviews of, and extended bibliographies concerning geometric flows, both analytically and numerically, and also including mean–field approaches.
A cylinder ${\cal C}(r)$ of radius $r$ is a solution to (4) iff
@@
However, $E,{\cal A}$ and $\cal V$} are not well defined for an infinite cylinder, and in the following we restrict to cylinders ${\cal C}_{L}(r)$ of lengths $L$. For these we have the scaling $E({\cal C}_{L}(\alpha r);c_{0})=E({\cal C}_{\alpha L}(r);c_{0}/\alpha)$ and thus it is sufficient to consider $r=1$ with a variable length $L$, and we abbreviate ${\cal C}_{L}={\cal C}_{L}(1)$. After showing well–posedness of (5) for initial conditions $X_{0}$ close to ${\cal C}_{L}$, where we assume periodic boundary conditions along the cylindrical axis, we study bifurcation diagrams of steady states of (5). In Fig.1(c) we preview four types of solutions from branches bifurcating from cylinders, obtained from numerical continuation and bifurcation, namely again pearling (A) and coiling (B) like in (a) and (b), and buckling (C), and wrinkling (D), which break the symmetry of the cylinder in different ways.  [^2] While pearling is a standard name for A, the nomenclature for B and in particular C varies. For instance, B is sometimes called writhing or twisting, and in [NSS15], C is called wrinkling. We could not find explicit names for states like D in the “vesicles literature”, and our choice of “wrinkling” is inspired by [FVKG22], which treats a related model for arteries with a restoring force.  [^3] The symmetry group of the problem at the cylinder is $O(2)\times O(2)$, translations and reflections in $x$, and in $\varphi$, where $x$ is along the cylindrical axis and $\varphi$ the polar angle orthogonal to it. Additionally there are the trivial symmetries of translation in $y$ and $z$, the euclidian coordinates orthogonal to $x$, but these are mostly relevant from a numerical point of view, i.e., must be dealt with via phase conditions in the numerical continuation. The starting point for the bifurcation analysis is the dispersion relation $\mu=\mu(m,n)$ for the linearization of (5) at ${\cal C}_{L}$, which can be computed explicitly from an ansatz
$X={\cal C}_{L}+{\rm e}^{\mu t+{\rm i}(k_{m}x+n\varphi)}N$,   $m,n\in{\mathbb{Z}}$,    $k_{m}=2\pi m/L$.
We study these dispersion relations with $c_{0}$ and $\lambda_{2}$ (and $\lambda_{1}$ given by (7)) as primary parameters, and for suitable $c_{0}$ and $L$ we find (typically narrow) ranges in $\lambda_{2}$ of stability of the “trivial branch” $X={\cal C}_{L}$, with branch points at the ends. At one end, axisymmetric branches of pearling type A bifurcate, while at the other end the loss of stability is to B,C or D, depending on the parameters.
After identifying these instabilities, based on the well–posedness results for the flow we use a centre manifold reduction to derive amplitude equations (AEs) for the critical modes, and from these also obtain approximations of the bifurcating steady state branches. The derivation of these AEs exploits the $O(2)\times O(2)$ symmetry of the problem, but is rather non–standard due to the constraints (2), and in some sense our methods could be described as bifurcation without parameters [Lie15], see Remark 1.
Finally, we use the package pde2path [Uec21, MU24b] to numerically continue the primary bifurcating branches away from their bifurcation points from ${\cal C}_{L}$. Close to bifurcation these numerical results agree well with the AE description, and additionally we obtain the behavior and stability at larger amplitude, and secondary bifurcations. In summary, this yields rather complicated bifurcation diagrams (BDs), which we discuss in detail for three exemplary cases, namely $c_{0}=0,c_{0}=-1$, and $c_{0}=0.48$, with strong qualitative differences, and we complement these steady state BDs with numerical flows for (5) to give further indication of the typical behavior of (5).
Remark 1.1.
We start the linear stability analysis and also the numerical bifurcation analysis by continuation of the trivial solution (branch) ${\cal C}_{L}$ in the Lagrange multiplier $\lambda_{2}$ (with $\lambda_{1}$ given by (7)). We then also often continue the nontrivial branches (e.g., of type A–D from Fig.1) in $\lambda_{2}$ as a primary parameter, with fixed $\cal A$} but without fixing $\cal V$}. However, the stability is always computed wrt to (5), i.e., for the area and volume conserving flow. This point of view of allowing the volume $\cal V$} to change along the branch while considering stability under $\cal V$} conservation is motivated by the closed vesicle situation discussed for example in [JSL93] and [MU24b]. Roughly speaking, for closed vesicles, the sphere is the unique maximum of volume at fixed area. Hence volume and area preserving branches bifurcating from the sphere do not exist. The same holds for the cylinder if we also fix the length $L$. Cylinders can be destabilized preserving $\cal V$} and $\cal A$} when allowing $L$ to change, see, e.g., [ZcH89], but we found our method with fixed $L$ more convenient. $\rfloor$
Remark 1.2.
a) Following [BZM94], [GNPS96] considers the dynamics of pearling based on the Helfrich energy and treating the hydrodynamics in the tube via a lubrication approximation. This leads to a 6th order Cahn–Hilliard type equation for axially symmetric perturbations of the cylinder, with $\cal V$} conservation, and the critical wave numbers and growth rates of the pearling (including the propagation speed of pearling fronts) are carefully compared to [BZM94]. The coiling instability and the dynamics of coiling are discussed [FTGB+99, TGBKS01], and compared to the dynamics of pearling in [TSA+01]. For both, it is argued that polymers anchoring at lipid bilayers locally generate nonzero values for the spontaneous curvature $c_{0}$, and generate significant differences in effective area between inside and outside of the bilayer(s), which points at area difference elasticity (ADE) terms. In particular, in [TSA+01], measurements of diffusion/propagation speeds suggest that for local injection of polymers in the outer fluid, initially ADE terms (with bulk diffusion of polymers) give the right model, and spontaneous curvature terms (with surface diffusion of polymers) when reaching equilibrium, both with $\cal A$} and $\cal V$} conservation. On the other hand, [CR98] argues that in certain experiments $c_{0}\neq 0$ is a global quantity. Finally, an important feature of coiling is that it rather occurs for multilamellar tubes, and that in equilibrium the coiling is usually “tight”, as in Fig.1(b), and “stretched coils” as B in (c) only occur transiently. See also [Sta01] for an overview of these results.
Further experiments on coiling are discussed in [PQAEM09], with various hypothesis for the underlying mechanism, and [HEJ+14] discusses the role of the actin cortex (not considered in (5)) in suppressing pearling in actual cells. In [NSS15], tubular vesicles modeled by (1) with $c_{0}=0$ and $\cal A$} and $\cal V$} conservation, are subjected to external elongational and compressional flows, and numerical simulations are compared to experimental results. Depending on flow parameters (condensed to the capillary number), and the initial vesicle shape (including, e.g., the shape of the spherical caps at the ends of finite cylinders), pearling is observed for elongational flows, and buckling and coiling (called wrinkling in [NSS15]) for compressional flows. This is also complemented by linear stability analysis for infinite cylinders.
b) In our model (5) we do not consider any interaction with surrounding fluids; we treat the idealized infinitely long cylinders, or the idealization of a fixed spatial periodicity; and $c_{0}$ and the surface tension $\lambda_{1}$ and pressure difference $\lambda_{2}$ are always global. Hence, mismatches to experiments (small range of stability for periodic pearling, long wave nature of the buckling and coiling, see below) are not completely surprising, but we believe/hope that our analytical (instabilities of ${\cal C}_{L}$, derivation of the AEs) and numerical (continuation of the bifurcating branches to larger amplitude) discussion of (5) will also help the further modeling. $\rfloor$
Outline, and overview of results
In §2 we recall the needed differential geometry and rewrite (5) as an equation in the normal displacement $u$ from the cylinder. This gives a constrained quasilinear parabolic problem in $u$, for which we use maximal regularity in Sobolev spaces to obtain local existence close to ${\cal C}_{L}$. From this we also obtain that nonlinear stability follows from linearized stability, and since ${\cal C}_{L}$ has constant mean and Gauss curvatures, the eigenvalues and eigenfunctions of the linearized problem can be computed explicitly allowing an analytical treatment of the instabilities of ${\cal C}_{L}$, and of the bifurcating branches via amplitude equations in §3 and §4. This is straightforward, but cumbersome due to the three parameters $c_{0},L$ and $\lambda_{2}$, and requires a careful look at the area and volume constraints (2) already in the linearization. The results of the linear stability and AE analysis can be summarized as follows:
  1. 1.
    The coiling/buckling instability of ${\cal C}_{L}$ is always of long wave type, i.e., the primary coiling/buckling instability has the maximal period given by the length $L$. This contradicts many experimental results for coiling, where the instability seems of finite wavelength type, and after onset the cylinders coils up tightly as in Fig.1(b). We could not find clear experimental results on buckling of vesicles.
  2. 2.
    For some (positive) $c_{0}$ the primary pearling instability can be of finite wave number type, i.e., independent of $L$, corresponding to Fig.1(a). However, see 5. below.
  3. 3.
    The wrinkling instability (homogeneous along the cylindrical axis) naturally is independent of $L$.
  4. 4.
    Due to the constraints, also the derivation and interpretation of the AEs is non–standard. In particular, the pearling and wrinkling branches bifurcating at stability loss of ${\cal C}_{L}$ are always stable. For the coiling vs buckling at stability loss of ${\cal C}_{L}$, always exactly one of these bifurcating branches is stable, and the instability of the unstable one is wrt perturbations into direction of the stable one.
In §5 we complement our analysis with numerical continuation and bifurcation methods, and also with some numerical flows for (5), aka direct numerical simulation (DNS). As expected and already said, near bifurcation the numerics agree well with the analysis, but one remarkable result is the following:
  1. 5.
    In case of finite wave number pearling the stability range of the bifurcating pearling branch rather strongly depends on $L$: The pearling branch becomes unstable via exchange of stability with a bifurcating branch of a “localized single pearl”, and the associated bifurcation point moves closer to the primary bifurcation with increasing $L$. This suggests that finite wavelength $O(1)$ amplitude pearling as in Fig.1(a) does not arise via bifurcations at stability loss of ${\cal C}_{L}$ in (5).
See §5 for details of this, and also for secondary bifurcations from other primary branches. On the other hand:
  1. 6.
    For the case of “large” $c_{0}$ ($c_{0}>1/2$, where ${\cal C}_{L}$ is not stable for any $(\lambda_{1},\lambda_{2})$) we do find finite wavelength $O(1)$ pearling, but not via bifurcation from ${\cal C}_{L}$. Instead we use DNS to converge to such solutions, which we subsequently can continue again in parameters. However, we only present this as an outlook in §5.4, and leave a systematic study of this setting for future work.
Another interesting result from the numerical flows is that cylindrical shapes can be more stable than linearized stability analysis of steady solutions $(X,\lambda_{1},\lambda_{2})$ at fixed ${\cal A}_{0}$ and ${\cal V}_{0}$ suggests. Namely:
  1. 7.
    Non area and/or volume conserving perturbations, of, e.g., unstable coiling, may yield initial conditions which converge to nearby stable coiling at the new conserved $({\cal A}_{0},{\cal V}_{0})$ by mainly adjusting the surface tension $\lambda_{1}(t)$ and the pressure $\lambda_{2}(t)$, with rather negligible changes of $X(t)$.
In §6 we summarize our analysis again, and in the Appendix we collect some details of the derivation of the AEs, and some remarks on the numerical algorithms.
@@ 2 Preliminaries and well-posedness of the flow
We introduce the basic differential geometric preliminaries in order to write (5) as a constrained parabolic quasi-linear equation in the normal displacement $u$ of the cylinder. This construction is as in [LS16], and hence we refrain from extensive computational details and focus on the structure and mapping properties of the resulting equations. From these we obtain maximal regularity results and hence local existence for (5). For general background on differential geometry we recommend [Tap16, UY17].
@@ 2.1 Preliminaries
$\begin{array}[]{rl}\Phi:{\mathbb{R}}\times\mathbb{T}&\rightarrow{\mathbb{R}}^{3}\\ (x,\varphi)&\mapsto\begin{pmatrix}x,r\cos(\varphi),r\sin(\varphi)\end{pmatrix}^{T}\end{array}$ (10)
of the product manifold ${\mathbb{R}}\times\mathbb{T}$, where $\mathbb{T}$ is the one dimensional topological torus with length $2\pi$ endowed with the metric $d_{\mathbb{T}}(\varphi,\vartheta)=\min\{|\varphi-\vartheta|,2\pi-|\varphi-\vartheta|\}$ for all $\varphi,\vartheta\in\mathbb{T}$. Then as the pullback metric of the ${\mathbb{R}}^{3}$ to ${\cal C}(r)$ we have $g_{0}=\,{\rm d}x^{2}+r^{2}\,{\rm d}\varphi^{2}$. The normal vector (field) $N$ of ${\cal C}(r)$ is oriented inwards, wrt $\Phi$, hence $H({\cal C}_{r})=\frac{1}{2r}$.
@@
assuming $u<r$, such that $X$ is still an embedding of ${\mathbb{R}}\times\mathbb{T}$; explicitly,
@@
Lemma 2.1.
For $u,v\in C^{\infty}_{\text{per}}({\cal C}_{L}(r))$ sufficiently small, the area of $X(u)={\cal C}_{L}(r)-u\,N$ expands as ${\cal A}(u+v)={\cal A}(u)+\left\langle{\partial}_{u}{\cal A}(u),v\right\rangle+\frac{1}{2}\left\langle{\partial}_{u}^{2}{\cal A}(u)v,v\right\rangle+O(\|v\|^{3})$ with
$\left\langle{\partial}_{u}{\cal A}(u),v\right\rangle{=}2\int_{X}H(u)\eta(u)v\,{\rm d}S,\ \ \left\langle{\partial}_{u}^{2}{\cal A}(u)v,v\right\rangle{=}\int_{X}-(\Delta_{u}v+2(H^{2}(u)-K(u))v)v\,{\rm d}S.$ (16)
The volume
@@
expands as ${\cal V}(u+v)={\cal V}(u)+\left\langle{\partial}_{u}{\cal V}(u),v\right\rangle+\frac{1}{2}\left\langle{\partial}_{u}^{2}{\cal V}(u)v,v\right\rangle+O(\|v\|^{3})$ with
$\left\langle{\partial}_{u}{\cal V}(u),v\right\rangle=\int_{X}\eta(u)v\,{\rm d}S,\qquad\left\langle{\partial}_{u}^{2}{\cal V}(u)v,v\right\rangle=\int_{X}g^{-1/2}(u)v^{2}\,{\rm d}S.$ (18)
Remark 2.2.
Since $\Lambda$ appears linearly in $G(u,\Lambda)$, the first derivative $\frac{\rm d}{{\rm d}t}q(u)=0$ gives a linear system of equations
$\mathcal{D}\Lambda=\begin{pmatrix}\left\langle{\partial}_{u}E,{\partial}_{u}{\cal A}(u)\right\rangle\\ \left\langle{\partial}_{u}E,{\partial}_{u}{\cal V}(u)\right\rangle\end{pmatrix}$, (19)
@@
$\mathcal{D}:=\left\langle{\partial}_{u}q(u),{\partial}_{\Lambda}G(u,\lambda)\right\rangle=\begin{pmatrix}4\int_{X}H^{2}(u)\eta^{2}(u)\,{\rm d}S&2\int_{X}H\eta^{2}(u)\,{\rm d}S\\ 2\int_{X}H\eta^{2}(u)\,{\rm d}S&\int_{X}\eta^{2}(u)\,{\rm d}S\end{pmatrix}$. (20)
If $\det\mathcal{D}=4\left(\int_{X}\eta^{2}\,{\rm d}S\int_{X}H^{2}\eta^{2}\,{\rm d}S-\left(\int_{X}H\eta^{2}\,{\rm d}S\right)^{2}\right)\neq 0$, then the implicit function theorem gives that in a neighborhood of $u$ there is a unique solution $\Lambda=\Lambda(u)$. Following [NY12], inserting $\Lambda(u)$ in (15a) gives a parabolic integro differential equation. However, if $u=0$, then $\det{\cal D}=0$, but the right hand side of (19) is orthogonal to $\ker{\cal D}$ (more generally, this holds at any constant mean curvature surface). Then a solution $\Lambda$ to (19) exists, and $\left\langle\Lambda_{\cal C},\ker{\cal D}^{T}\right\rangle=0$ gives that $2\lambda_{1}=\lambda_{2}$ and $\lambda_{1}=\frac{1}{3{\cal A}_{0}}\left\langle{\partial}_{u}E,1\right\rangle$, i.e., $\ker{\cal D}$ corresponds to the family of cylinders in (7); see [NY12] for further discussion (for the case of closed vesicles of spherical topology).